A particle moving with a uniform acceleration along a straight line covers distances a and b in successive intervals of p and q second. The acceleration of the particle is

Grade:12

1 Answers

dara shannu
9 years ago
diatance travelled by particle in p th second = u+A(2p-1)/2 = a
where u is initial velocity and A is acceleration
and in q th second =u+A(2q-1)/2 = b
solving both we get
                                   A = a – b/p – q